OpenStudy (anonymous):

One a scale of 1/2" = 1', what would be the length of a line representing 10'3"?

OpenStudy (anonymous):

Five feet, one and one-half inches.

OpenStudy (radar):

The 10'3" would be 10 1/4 ft, multiply that by 1/2 " =(10.25)(.5)=5.125 or the line would be 5 1/8 in. long.

OpenStudy (anonymous):

I don't quite understand your question, but here's my best guess. 8" or 11'1" 1/2" = 1' --> 10' = 5" 5" + 3" = 8" or 10'3" 3" = 1.5' = 11.5' or 11'1"

OpenStudy (radar):

We need a draftsman lol

OpenStudy (anonymous):

In the answer, 5.125, the decimal portion indicates 1/8 of a foot, which is 1.5 inches.

OpenStudy (anonymous):

The answer is 5 1/4, but I don't know how to get to it!

OpenStudy (radar):

Look at it like this The 10' 3" is equal to 10' plus .25 ft which is the 3 inches. Now our scale is 1/2" is equal to 1 ft. Right off we know the 10' would be represented by 5 inches (as their is (10) (1/2) in 5 That leaves the 3 in or .25 feet if 1/2 in represents 1 foot, then 1/4 in is 1/2 ft then 1/8 in equal 1/4 feet. again I say 5 1/8 in or 5.125 in. represents the required 10' 3 " dimension

OpenStudy (anonymous):

what was the scale part all about ...?

OpenStudy (anonymous):

An architects scale...

OpenStudy (anonymous):

like for plans.

OpenStudy (anonymous):

Convert the 10'3' from inches and feet to feet...10.25 feet Multiply 10.25 feet times 1/2 (or .5) to get 5.125 FEET 5.125 feet, when converted to feet and inches, is 5 feet, 1 and 1/2 inches.

OpenStudy (radar):

To verify lets see if 5 1/4 divided by 1/2 meets the required 10' 3"

OpenStudy (radar):

5.25/.5 = 10.5 feet or 10 ft 6 in Not 10' 3"

OpenStudy (radar):

Again the answer if the problem is stated correctly is 5.125 in.

OpenStudy (anonymous):

5.125 inches??? One half of 10 feet alone is 5 feet...the answer cannot be 5.125 inches!

OpenStudy (radar):

Remember rebeccablu that 1/2 in = 1 ft on the stated scale. So why wouldn't a 5.125 in line represent a 10' 3" dimension using that scale. Inquiring minds need to know!

OpenStudy (anonymous):

radar! i now see what i did not see...that one/half INCH equals one foot. mea culpa! My figures were bases on one/half foot equals one foot. in the words of gilda radner's character, emily latella,,, never mind!

OpenStudy (anonymous):

I think it's stated correctly. this is coming from a practice test I'm taking.

OpenStudy (anonymous):

I have one more...same idea if you have another minute to help me out!

OpenStudy (radar):

OH I'm so glad that you found the error of your thinking, I knew you were convinced you were right, and I so wanted to get you to thinking about this problem. No Problem now does lindsy9 see how it was solved?

OpenStudy (anonymous):

On a scale of 1/8"=1', 3'3" would be represented by by what fraction of an inch?

OpenStudy (radar):

do you want it rebeccablu?

OpenStudy (radar):

lindsay while rebeccablu is thinking are you sure the answer said 5.25 or did it say 5.125???

OpenStudy (anonymous):

it says 5 1/4"...but maybe they are wrong!

OpenStudy (radar):

O.K to your last problem the scale is 1/8 " equals 1 ft. Right away we know 1 in is 8 ft, so for sure 3' 3" is less than a inch. First convert the 3' 3" dimension to ft and that is 3.25 (3" is 1/4 of a ft or .25 ft.) next divide the 3.25 by 1/8 and since your answer is wanted to be in fractions i will convert to fractions 3 1/4 is equal to improper fraction of 13/4 13 1 13 --- X ----- = ----- in. 4 8 32

OpenStudy (radar):

Sorry change divide to multply made a typo lol,it is not division!

OpenStudy (radar):

Answer is still 13/32 (not even a half inch)

OpenStudy (anonymous):

Daaang, yep! That's the correct answer! Thanks so much for explaining it to me! I will study this, understaND it and hopefully get it right on the test!

OpenStudy (radar):

Hey good luck with that test, noticed I multiplied not divided made an error said one thing but did another, what I did was right multiply. To verify answer you would divide by 1/8.

OpenStudy (radar):

Don't know why the first one did not agree with the "approved solution" was the length 10' 3" ?

OpenStudy (radar):

If it was 10' 6" that would be correct just thought you copied it down wrong.